# Average Rényi Entanglement Entropy in Gaussian Boson Sampling

### v1

@@ 1 Introduction
There is strong evidence that classically sampling from the output probability distribution of a GBS experiment is hard in the complexity-theoretic sense [7, 8, 9, 10, 11], whereas such sampling can be done efficiently on a quantum device. Moreover, entanglement is a critical asset across quantum computing, with additional applications in teleportation, dense coding, and quantum communication [12, 13, 14, 15]. Therefore, quantifying the behavior of entanglement in Gaussian bosonic systems has been of recent interest [7, 8], especially to sharpen the connection between entanglement and computational complexity.
@@
In 1993, Page conjectured an explicit formula for the average entanglement entropy over all Haar-random pure states in an $N$-dimensional Hilbert space of the reduced density matrix on an $rN$-dimensional subsystem for $r\in[0,1]$. This conjecture was proven soon after in the $N\to\infty$ limit, and the average entanglement entropy as a function of $r$ is now known as the Page curve [16, 17, 18]. The Page curve was initially studied in the context of information dynamics in and around black holes [19, 20, 21, 22], and has since found applications in open quantum systems [23, 24] and quantum information theory [25, 26, 27]. As a natural follow-up, many works investigated the deviation of entanglement entropy from its expected value. A system exhibits typical entanglement if the probability that a random state has entanglement bounded away from the average is small. This behavior, along with strong and weak classes of entanglement typicality, has been studied in black holes, thermodynamics, and other systems [28, 29, 30, 31, 32].
@@
Leaving the realm of standard finite-dimensional systems, the Page curve for fermionic Gaussian states has been extensively analyzed [33, 34, 35]. A natural next step, especially given the relevance to GBS, is to study bosonic Gaussian states. The typicality of entanglement in random bosonic Gaussian states was studied in Refs. [36, 37, 38, 39], and the Page curve for Gaussian bosonic systems was first studied in Ref. [39]. However, Ref. [39] only computed the Page curve for the Rényi-2 entropy.
@@
In this work, we derive analytic expressions for the Rényi-$\alpha$ Page curves of bosonic Gaussian systems for all $\alpha\in\mathbb{Z}_{>0}$, with special considerations for the von Neumann (Rényi-1) Page curve. The entropy of a mixed state is an indicator of its entanglement with the rest of the universe [40], and indeed the Rényi-$\alpha$ Page curve is a measure of the entanglement of an average state in a system as measured by the Rényi-$\alpha$ entropy, for $\alpha\in\mathbb{R}^{+}$. Each Rényi-$\alpha$ entropy has its own unique mathematical features, and notable examples are the von Neumann/Shannon entropy ($\alpha=1$), collision entropy ($\alpha=2$), and min-entropy ($\alpha\to\infty$). In particular, the min-entropy, the entropy defined by $-\log\text{max}({p_{i}})$ for each $p_{i}$ a probability of occurence of outcome $i$ when working with a discrete random variable, is a useful expression for working with extracting randomness from variables when lower entropies (e.g. von Neumann) do not suffice [41]. Therefore, it is natural to analyze the Rényi-$\alpha$ entropies in bosonic systems and extract their unique properties.
@@
Our expressions are asymptotically exact in the number of modes $n$. Using these expressions, we prove various results about typicality of entanglement with respect to these Rényi-$\alpha$ entropies. Our setup is exactly that of a GBS experiment, and is the same as the setups considered in Refs. [38, 39]. A pure input quantum state with $n$ squeezed vacuum modes is acted on by a Haar-random, linear unitary (passive Gaussian) circuit. The Page curve, as a function of $r$, is the average Rényi-$\alpha$ entanglement between a group of $rn$ modes and the remaining system of $(1-r)n$ modes, where $r\in[0,1].$ We expound our setup in greater detail in Section 2.
@@
In Section 3, we present our results for the von Neumann entropy. Under an equal squeezing regime, we first present an explicit formula for the von Neumann Page curve in Section 3.1 in terms of the partition ratio $r$ and initial squeezing strength $s$, which is exact up to terms that decay to zero as $n\rightarrow\infty.$ We subsequently discuss the von Neumann entropy under squeezing limits $s\rightarrow 0$ and $s\rightarrow\infty$ in Section 3.2.
@@
We then generalize our analyses to the Rényi-$\alpha$ entropies for integers $\alpha\geq 2$ in Section 4. In Section 4.1, we derive the Rényi-$\alpha$ Page curve in terms of the partition ratio $r$ and initial squeezing strength $s$, with additional terms that vanish to zero as $n\rightarrow\infty$. We calculate squeezing limits of the Rényi-$\alpha$ entropy, this time with the extra dependence on $\alpha$, in Section 4.2. Finally, we evaluate the Rényi-$\alpha$ entropy under unequal, small initial squeezing strengths in Section 4.3 and derive typicality of entanglement results in Section 4.4.
2 Setup
@@ 3.2 Squeezing limits
While the derivation of Theorem 2 is fairly straightforward, its significance is more interesting. The von Neumann Page curve’s growth in order $s^{2}\log(1/s^{2})$ for small $s$ can be attributed to the unique definition of the von Neumann entropy $S_{1}(\rho)=-\operatorname{Tr}(\rho\log\rho),$ where $\rho$ is the density matrix of the quantum state. Therefore, the symplectic eigenvalues $\nu_{j}$ of the covariance matrix would also scale the von Neumann entropy in order $\nu_{j}\log\nu_{j},$ warranting the $s^{2}\log(1/s^{2})$ dependence (as $\nu_{j}\sim s^{2}$). Furthermore, the von Neumann entropy is non-analytic unlike the rest of the Rényi entropies. Therefore, the argument presented in Ref. [39] that the Rényi entropy can be expanded in lowest order $s^{2}$ doesn’t apply to the von Neumann case. More specifically, the deviation from analytic behavior arises from the von Neumman entropy formula in terms of the covariance matrix $\sigma$ [47, 42, 48]:
@@
Here, the von Neumann Page curve scales with $s$ as $s\rightarrow\infty.$ This is because, for large absolute squeezing values, higher orders in $r$ dominate the expansion of Theorem 1, which scale with $s.$ Interestingly, Theorem 3 remains the exact same for the Rényi-2 entropy [39], and indeed we will see in Sec. 4.2 that it is the same for all the Rényi-$\alpha$ entropies for integral $\alpha$.
3.3 Unequal squeezing
@@
4 General Rényi-$\alpha$ Entropies
[…]
@@ 4 General Rényi-$\alpha$ Entropies
In this section, we extend much of our analysis with the von Neumann entropy in Section 3 to Rényi-$\alpha$ entropies for integer $\alpha\geq 2$. We also derive results for the typicality of entanglement and the small, unequal squeezing case. First, we present the Rényi-$\alpha$ Page curve $\forall\alpha\in\mathbb{Z},\alpha\geq 2.$ Next, we investigate squeezing limits as $s\rightarrow 0$ and $s\rightarrow\infty,$ as well as consider the Rényi-$\alpha$ Page curve for small, unequal squeezing strengths. We also study the variance of the entanglement and prove various regimes of entanglement typicality. Finally, we show some numerical simulations confirming our results.
4.1 Explicit formula
@@
$\begin{split}\underset{U\in U(n)}{\mathbb{E}}S_{\alpha}(U)=\frac{\zeta}{\alpha-1}\cdot\underset{U\in U(n)}{\mathbb{E}}S_{2}+\frac{1}{\alpha-1}\sum_{i=1}^{\infty}\left[\sum_{m=1}^{a}\frac{\sinh^{2i}(2s)}{i(\cosh^{2}(2s)+\cot^{2}\frac{\pi m}{\alpha})^{i}}\right]\left(nG_{i}(r)-H_{i}(r)+o(1)\right),\end{split}$ (8)
@@
Since $G_{i}(r)$ is symmetric under $r\to 1-r$, we see that the Page curves are also manifeslty symmetric under $r\to 1-r$. Various analytic Rényi-$\alpha$ Page curves are plotted in Fig. 1. We also plot numerically-simulated values of the Page curve, where we generate Haar-random unitary matrices and compute the entropies using the positive symplectic eigenvalues $\nu_{j}$ of each covariance matrix $\sigma$. Then, we compute the respective Rényi-$\alpha$ entropy as $S_{\alpha}(U)=\sum_{j=1}^{n}\frac{1}{1-\alpha}\ln\left(\frac{2^{\alpha}}{(\nu_{j}+1)^{\alpha}-(\nu_{j}-1)^{\alpha}}\right)$ [42]. The numerical and analytical values overlap, indicating the validity of Theorem 1 and Theorem 4.
4.2 Squeezing limits
@@
In the infinite squeezing case, the Page curve remains independent of $\alpha$ in order $s$ as $\alpha\rightarrow\infty$. This indicates that, as $s\to\infty,$ the Page curve approaches the curve of maximum entanglement.
$k=\Theta\!\left(n\right)$ $k=o\!\left(n\right)$ [39] $k=o\!\left(n^{1/3}\right)$ [38]
@@
Table 1: A summary of the current status of rigorous results on entanglement typicality in Gaussian bosonic systems. This table is the same as the one provided in Ref. [39], but the Rényi-2 case has been generalized for all Rényi-$\alpha$ entropies where $\alpha\in\mathbb{Z}^{+}.$ Note that “weak” indicates that the result is not fully proven, but is based on Conjecture 10 in Ref. [39]. Furthermore, wherever "weak" typicality is displayed, we have not ruled out the possibilty of strong typicality. We present typicality results when $k,$ the number of modes in the subsystem, scales with $\Theta\!\left(n\right),o\!\left(n\right),$ and $o\!\left(n^{1/3}\right),$ respectively. Note that we only take the "worst-case" typicality; if there exists both weak and strong typicality for one scaling, we display only the weakly typical case. We present typicality results separately for the Rényi-$\alpha$ entropies and the von Neumann entropy. The leftmost column ($k=\Theta\!\left(n\right)$) stems from our findings in Section 4.4, while the middle column ($k=o\!\left(n\right)$) follows from Ref. [39]. The rightmost column ($k=o\!\left(n^{1/3}\right)$) follows from the results of Ref. [38]. Refs. [36, 37] proved strong typicality in the regime $k=\mathcal{O}\!\left(1\right)$.
4.3 Unequal squeezing
@@ 4.4 Typicality of entanglement
5 Conclusions and open questions
@@
Another problem to tackle deals with unequal squeezing. There are two primary avenues of future research. First, for small, varying squeezing strengths $s_{1},s_{2},\dots,s_{n},$ what is the average von Neumann entanglement entropy of the subsystem? We have outlined a method in Section 3.3 of tackling this problem, but initial attempts have proved futile due to the heavy presence of inverse matrices. Nonetheless, learning the behavior of the von Neumann entropy for unequal squeezing regimes is important from a practical standpoint, since many GBS experiments are conducted with one set of modes at zero squeezing strength and the other at some varied squeezing strength close to zero. Second, it is interesting to consider Rényi-$\alpha$ entropies for systems with varying squeezing strengths that are not necessarily close to zero. The argument given in Section 4.3 revolves around the entropy expression being analytic around $s=0$ and vanishing in higher orders of $s_{\text{max}}.$ However, for the most general case of $s_{\text{max}}\in\mathbb{R},$ one cannot simply extend the $s\rightarrow 0$ squeezing limit for unequal squeezing. Nonetheless, deriving such an expression would be important for a wider range of bosonic systems.
Acknowledgements
This work was supported in part by the DoE ASCR Accelerated Research in Quantum Computing program (award No. DE-SC0020312), DARPA SAVaNT ADVENT, AFOSR MURI, DoE ASCR Quantum Testbed Pathfinder program (awards No. DE-SC0019040 and No. DE-SC0024220), NSF QLCI (award No. OMA-2120757), NSF STAQ program, and AFOSR. Support is also acknowledged from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator. JTI thanks the Joint Quantum Institute at the University of Maryland for support through a JQI fellowship. Specific product citations are for the purpose of clarification only, and are not an endorsement by the authors or NIST.
[…]

### v2

@@ I Introduction
There is strong evidence that classically sampling from the output probability distribution of a GBS experiment is hard in the complexity-theoretic sense [7, 8, 9, 10, 11], whereas such sampling can be done efficiently on a quantum device. Moreover, entanglement is a critical asset across quantum computing, with additional applications in teleportation, dense coding, and quantum communication [12, 13, 14, 15]. Therefore, quantifying the behavior of entanglement in Gaussian bosonic systems has been of recent interest [7, 8], especially to sharpen the connection between entanglement and computational complexity. Indeed, the relationship between entanglement and complexity is generally unknown. Classical hardness requires at least some entanglement, but too much entanglement can also cause a state to be useless for computation. In fact, a randomly chosen finite-dimensional quantum state generally has too much entanglement to be computationally useful [16, 17]. This prompts our study of average and typical entanglement, and motivates further study into the general connection between entanglement and complexity.
@@
In 1993, Page conjectured an explicit formula for the average entanglement entropy over all Haar-random pure states in an $N$-dimensional Hilbert space of the reduced density matrix on an $rN$-dimensional subsystem for $r\in[0,1]$. This conjecture was proven soon after in the $N\to\infty$ limit, and the average entanglement entropy as a function of $r$ is now known as the Page curve [18, 19, 20]. The Page curve was initially studied in the context of information dynamics in and around black holes [21], and has since found various applications in quantum information and condensed matter theory [22]. As a natural follow-up, many works investigated the deviation of entanglement entropy from its expected value. A system exhibits typical entanglement if the probability that a random state has entanglement bounded away from the average is small. This behavior, along with strong and weak classes of entanglement typicality, has been studied in black holes, thermodynamics, and other systems [22, 23, 24, 25, 26, 27, 28].
@@
Leaving the realm of standard finite-dimensional systems, the Page curve for fermionic Gaussian states has been extensively analyzed [29, 30, 31, 32, 33, 34, 35]. A natural next step, especially given the relevance to GBS, is to study bosonic Gaussian states. The typicality of entanglement in random bosonic Gaussian states was studied in Refs. [36, 37, 38, 39], and the Page curve for Gaussian bosonic systems was studied in Ref. [39]. However, Ref. [39] only computed the Page curve for the Rényi-2 entropy.
@@
In this work, we derive analytic expressions for the Rényi-$\alpha$ Page curves of bosonic Gaussian systems for all $\alpha\in\mathbb{Z}_{>0}$, with special considerations for the von Neumann (Rényi-1) Page curve. The entropy of a mixed state is an indicator of its entanglement with the rest of the universe [40], and indeed the Rényi-$\alpha$ Page curve is a measure of the entanglement of an average state in a system as measured by the Rényi-$\alpha$ entropy, for $\alpha\in\mathbb{R}^{+}$. The Rényi-$\alpha$ entropies offer a quantitative way to compare the amount of entanglement in two given quantum states, in that if the Rényi-$\alpha$ entanglement entropy of state $\mathinner{\lvert\psi_{1}\rangle}$ is greater than $\mathinner{\lvert\psi_{2}\rangle}$ for any given $\alpha$, then one cannot convert $\mathinner{\lvert\psi_{2}\rangle}$ into $\mathinner{\lvert\psi_{1}\rangle}$ via local operations and classical communication. Each Rényi-$\alpha$ entropy has its own unique mathematical features, and notable examples are the von Neumann/Shannon entropy ($\alpha=1$), collision entropy ($\alpha=2$), and min-entropy ($\alpha\to\infty$). In particular, the min-entropy, the entropy defined by $-\log\text{max}({p_{i}})$ for each $p_{i}$ a probability of occurrence of outcome $i$ when working with a discrete random variable, is a useful expression for working with extracting randomness from variables when lower entropies (e.g. von Neumann) do not suffice [41]. Therefore, it is natural to analyze the Rényi-$\alpha$ entropies in bosonic systems and extract their unique properties.
Our expressions are asymptotically exact in the number of modes $n$. Using these expressions, we prove various results about typicality of entanglement with respect to these Rényi-$\alpha$ entropies. A surprising feature that our analysis reveals is that Rényi-$\alpha$ entropies all concentrate for integer $\alpha$, which suggests that the entanglement spectrum of the states generated by Gaussian boson sampling also concentrate. This result could be useful in theoretical analysis of the performance of Gaussian boson sampling circuits.
@@
Our setup is exactly that of a GBS experiment, and is the same as the setups considered in Refs. [38, 39]. A pure input quantum state with $n$ squeezed vacuum modes is acted on by a Haar-random, linear unitary (passive Gaussian) circuit. The Page curve, as a function of $r$, is the average Rényi-$\alpha$ entanglement between a group of $rn$ modes and the remaining system of $(1-r)n$ modes, where $r\in[0,1].$ We expound our setup in greater detail in Section II.
@@
In Section III, we present our results for the von Neumann entropy. Under an equal squeezing regime, we first present an explicit formula for the von Neumann Page curve in Section III.1 in terms of the partition ratio $r$ and initial squeezing strength $s$, which is exact up to terms that decay to zero as $n\rightarrow\infty.$ We subsequently discuss the von Neumann entropy under squeezing limits $s\rightarrow 0$ and $s\rightarrow\infty$ in Section III.2.
@@
We then generalize our analyses to the Rényi-$\alpha$ entropies for integers $\alpha\geq 2$ in Section IV. In Section IV.1, we derive the Rényi-$\alpha$ Page curve in terms of the partition ratio $r$ and initial squeezing strength $s$, with additional terms that vanish to zero as $n\rightarrow\infty$. We calculate squeezing limits of the Rényi-$\alpha$ entropy, this time with the extra dependence on $\alpha$, in Section IV.2. Finally, we evaluate the Rényi-$\alpha$ entropy under unequal, small initial squeezing strengths in Section IV.3 and derive typicality of entanglement results in Section IV.4.
II Setup
@@ III.2 Squeezing limits
We note that of course the entanglement must be zero when $s=0$ simply because there are no bosons present when $s=0$. Theorem 2 addresses how the entanglement approaches zero as $s\to 0$. While the derivation of Theorem 2 is fairly straightforward, its significance is more interesting. The von Neumann Page curve’s growth in order $s^{2}\log(1/s^{2})$ for small $s$ can be attributed to the unique definition of the von Neumann entropy $S_{1}(\rho)=-\operatorname{Tr}(\rho\log\rho),$ where $\rho$ is the density matrix of the quantum state. Therefore, the symplectic eigenvalues $\nu_{j}$ of the covariance matrix would also scale the von Neumann entropy in order $\nu_{j}\log\nu_{j},$ warranting the $s^{2}\log(1/s^{2})$ dependence (as $\nu_{j}\sim s^{2}$). Furthermore, the von Neumann entropy is non-analytic unlike the rest of the Rényi entropies. Therefore, the argument presented in Ref. [39] that the Rényi entropy can be expanded in lowest order $s^{2}$ doesn’t apply to the von Neumann case. More specifically, the deviation from analytic behavior arises from the von Neumman entropy formula in terms of the covariance matrix $\sigma$ [46, 42, 47]:
@@
Here, the von Neumann Page curve scales with $s$ as $s\rightarrow\infty.$ This is because, for large absolute squeezing values, higher orders in $r$ dominate the expansion of Theorem 1, which scale with $s.$ Interestingly, Theorem 3 remains the exact same for the Rényi-2 entropy [39], and indeed we will see in Sec. IV.2 that it is the same for all the Rényi-$\alpha$ entropies for integral $\alpha$.
We now compare this result to the case of averaging over all $n$-mode bosonic states with a fixed boson number constraint. In particular, the volume (i.e. proportional to $n$) scaling of the corresponding Page curve is [48]
$\begin{split}\lim_{n\to\infty}&\frac{1}{n}S_{1}^{(n_{b})}=\\ &\left((r_{b}+1)\log(1+r_{b})-r_{b}\log(r_{b})\right)\min(r,1-r),\end{split}$ (8)
where $r_{b}n=n_{b}$ the number of bosons. In this case, the Hilbert space is finite-dimensional with dimension $\binom{n+n_{b}-1}{n_{b}}$. As $n_{b}\to\infty$, this yields $\frac{1}{n}S_{1}^{(n_{b})}\sim\log(r_{b})\min(r,1-r)$
In our Gaussian setup, the expected number of bosons is $\left\langle n_{b}\right\rangle=n\sinh^{2}(s)$. In this case, the Hilbert space in infinite-dimensional, but the fixed squeezing constraint gives an energy constraint. Using Theorem 3, we then see that as $n_{b}\to\infty$ , $\frac{1}{n}S_{1}\sim\log\!\left(\left\langle r_{b}\right\rangle\right)\min(r,1-r)$.
We therefore see that, asymptotically, the average entanglement scales proportional to the volume of the system. Given a fixed but asymptotically large energy (i.e. number of bosons), the average entanglement over Gaussian states equals the average entanglement over all bosonic states.
Finally, we also note that when $r\to 0$, all of the Page curves that we have computed, and indeed all of the Page curves computed in Ref. [48], scale proportionally to $r$. In particular, we can see from Ref. [48, Fig. 2, Eq. (23)] that the fixed particle number Page curves for (a) fermions, (b) hardcore bosons, and (c) (multi-species) bosons all have Page curves that are proportional to $r$ as $r\to 0$. Similarly, Page curves in finite-dimensional spaces without constraints are also proportional to $r$ as $r\to 0$ [18, 19, 20, 22, 23, 24, 25, 26, 27, 28]. In the Gaussian case that we have computed in this work, the Page curve is also proportional to $r$ in this limit, as can be seen in Theorems 1 and 4.
III.3 Unequal squeezing
@@
IV General Rényi-$\alpha$ Entropies
@@
In this section, we extend much of our analysis with the von Neumann entropy in Section III to Rényi-$\alpha$ entropies for integer $\alpha\geq 2$. We also derive results for the typicality of entanglement and the small, unequal squeezing case. First, we present the Rényi-$\alpha$ Page curve $\forall\alpha\in\mathbb{Z},\alpha\geq 2.$ Next, we investigate squeezing limits as $s\rightarrow 0$ and $s\rightarrow\infty,$ as well as consider the Rényi-$\alpha$ Page curve for small, unequal squeezing strengths. We also study the variance of the entanglement and prove various regimes of entanglement typicality. Finally, we show some numerical simulations confirming our results.
IV.1 Explicit formula
@@
$\begin{split}\underset{U\in U(n)}{\mathbb{E}}S_{\alpha}(U)=\frac{\zeta}{\alpha-1}~{}\underset{U\in U(n)}{\mathbb{E}}S_{2}+\frac{1}{\alpha-1}\sum_{i=1}^{\infty}\left[\sum_{m=1}^{a}\frac{\sinh^{2i}(2s)}{i(\cosh^{2}(2s)+\cot^{2}\frac{\pi m}{\alpha})^{i}}\right]\left(nG_{i}(r)-H_{i}(r)+o(1)\right),\end{split}$ (9)
@@
Since $G_{i}(r)$ is symmetric under $r\to 1-r$, we see that the Page curves are also manifestly symmetric under $r\to 1-r$. Various analytic Rényi-$\alpha$ Page curves are plotted in Fig. 1. We also plot numerically-simulated values of the Page curve, where we generate Haar-random unitary matrices and compute the entropies using the positive symplectic eigenvalues $\nu_{j}$ of each covariance matrix $\sigma$. Then, we compute the respective Rényi-$\alpha$ entropy as $S_{\alpha}(U)=\sum_{j=1}^{n}\frac{1}{1-\alpha}\ln\left(\frac{2^{\alpha}}{(\nu_{j}+1)^{\alpha}-(\nu_{j}-1)^{\alpha}}\right)$ [42]. The numerical and analytical values overlap, indicating the validity of Theorem 1 and Theorem 4.
IV.2 Squeezing limits
@@
In the infinite squeezing case, the Page curve remains independent of $\alpha$ in order $s$ as $\alpha\rightarrow\infty$. This indicates that, as $s\to\infty,$ the Page curve approaches the curve of maximum Gaussian entanglement (that is, the maximum is taken over all Gaussian states with a fixed initial squeezing configuration).
$k=\Theta\!\left(n\right)$ $k=o\!\left(n\right)$ [39] $k=o(n^{1/3})$ [38]
@@
Table 1: A summary of the current status of rigorous results on entanglement typicality in Gaussian bosonic systems. This table is the same as the one provided in Ref. [39], but the Rényi-2 case has been generalized for all Rényi-$\alpha$ entropies where $\alpha\in\mathbb{Z}^{+}.$ Note that “weak” indicates that the result is not fully proven, but is based on Conjecture 10 in Ref. [39]. Furthermore, wherever "weak" typicality is displayed, we have not ruled out the possibilty of strong typicality. We present typicality results when $k,$ the number of modes in the subsystem, scales with $\Theta\!\left(n\right),o\!\left(n\right),$ and $o\!\left(n^{1/3}\right),$ respectively. Note that we only take the "worst-case" typicality; if there exists both weak and strong typicality for one scaling, we display only the weakly typical case. We present typicality results separately for the Rényi-$\alpha$ entropies and the von Neumann entropy. The leftmost column ($k=\Theta\!\left(n\right)$) stems from our findings in Section IV.4, while the middle column ($k=o\!\left(n\right)$) follows from Ref. [39]. The rightmost column ($k=o\!\left(n^{1/3}\right)$) follows from the results of Ref. [38]. Refs. [36, 37] proved strong typicality in the regime $k=\mathcal{O}\!\left(1\right)$.
IV.3 Unequal squeezing
@@ IV.4 Typicality of entanglement
V Conclusions and open questions
@@
Another problem to tackle deals with unequal squeezing. There are two primary avenues of future research. First, for small, varying squeezing strengths $s_{1},s_{2},\dots,s_{n},$ what is the average von Neumann entanglement entropy of the subsystem? We have outlined a method in Section III.3 of tackling this problem, but initial attempts have proved futile due to the heavy presence of inverse matrices. Nonetheless, learning the behavior of the von Neumann entropy for unequal squeezing regimes is important from a practical standpoint, since many GBS experiments are conducted with one set of modes at zero squeezing strength and the other at some varied squeezing strength close to zero. Second, it is interesting to consider Rényi-$\alpha$ entropies for systems with varying squeezing strengths that are not necessarily close to zero. The argument given in Section IV.3 revolves around the entropy expression being analytic around $s=0$ and vanishing in higher orders of $s_{\text{max}}.$ However, for the most general case of $s_{\text{max}}\in\mathbb{R},$ one cannot simply extend the $s\rightarrow 0$ squeezing limit for unequal squeezing. Nonetheless, deriving such an expression would be important for a wider range of bosonic systems.
Acknowledgements.
This work was supported in part by the U.S. Department of Energy, Office of Science, Accelerated Research in Quantum Computing, Fundamental Algorithmic Research toward Quantum Utility (FAR-Qu). Additional support is acknowledged from DARPA SAVaNT ADVENT, AFOSR MURI, DoE ASCR Quantum Testbed Pathfinder program (awards No. DE-SC0019040 and No. DE-SC0024220), NSF QLCI (award No. OMA-2120757), NSF STAQ program, AFOSR, ARL (W911NF-24-2-0107), and NQVL:QSTD:Pilot:FTL Support is also acknowledged from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator. JTI thanks the Joint Quantum Institute at the University of Maryland for support through a JQI fellowship. Y.-X.W. acknowledges support from a QuICS Hartree Postdoctoral Fellowship. Specific product citations are for the purpose of clarification only, and are not an endorsement by the authors or NIST.